OpenStudy (anonymous):

write an algebraic expression that is equivalent to cos13pi/12 using related acute angles

OpenStudy (anonymous):

maybe try \(\frac{13\pi}{12}=\frac{3\pi}{4}=\frac{\pi}{3}\)

OpenStudy (anonymous):

oops typo sorry i meant \[\frac{13\pi}{12}=\frac{3\pi}{4}+\frac{\pi}{3}\]

OpenStudy (anonymous):

then you can use the addition angle formula

OpenStudy (anonymous):

hm.. i used the formula \[\cos(\pi + \theta)=-\cos \theta\] since if i convert \[\frac{ 13\pi }{ 12 }\] into degrees it would be in the third quadrant

OpenStudy (anonymous):

this would work right?

OpenStudy (anonymous):

forget degrees, it is in the third quadrant because \[\pi<\frac{13\pi}{12}<\frac{3\pi}{2}\]

OpenStudy (anonymous):

ah true alright.

OpenStudy (anonymous):

you can evaluate this because \[\cos(\frac{13\pi}{12})=\cos(\frac{3\pi}{4}+\frac{\pi}{3})\] \[=\cos(\frac{3\pi}{4})\cos(\frac{\pi}{3})-\sin(\frac{3\pi}{4})\sin(\frac{\pi}{3})\] and we know all of those

OpenStudy (anonymous):

oooooooh okay thank you!

OpenStudy (anonymous):

yw